OpenStudy (anonymous):

Simon invested $1,550 at 6.5% simple interest. He earned $302.25 in interest after t years. What is the value of t? 6 9 3 1

OpenStudy (anonymous):

Its C 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

6.5% of 1,550 = 1,550 * .065 = 100.75 per year 100.75 * T years = 302.25 T = 302.25/100.75 = 3 years answer B

OpenStudy (welshfella):

each year he earns $ 1550 * 6.5/100 work this out then divide it into 302.25 to get t.

OpenStudy (anonymous):

YW :)

OpenStudy (anonymous):

Suzanne deposited $2,500 in a savings account earning 3.5% simple interest. What was the total interest earned after 6 years? $4,500 $52.50 $450 $525

OpenStudy (welshfella):

$2500* 0.035*6

OpenStudy (welshfella):

grab ur calculator and work it out

OpenStudy (anonymous):

Its 525

OpenStudy (zpupster):

the simple interest formula is I = Prt i= Interest P=principal R= rate t = time just plug and solve for the missing piece

OpenStudy (anonymous):

thats umm 450 right?

OpenStudy (welshfella):

@addictbeau - please dont give direct answers Read the Code of Conduct for this site

OpenStudy (anonymous):

Sorry I was going to explain it afterwards :)

OpenStudy (anonymous):

is it 450?

OpenStudy (anonymous):

No its not

OpenStudy (anonymous):

oh 525?

OpenStudy (anonymous):

Yes bc if you do $2500*0.035*6 you get 525 :)

OpenStudy (anonymous):

o ok thank you

OpenStudy (anonymous):

No problem